OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 2. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (academicgurusinc):

log1632 = x + 2 The solution to a logarithm is the exponent. The base number of a log is 10. 10(x + 2) = 1632 Log both sides of the equation and bring the exponent down as the coefficient of the log. (x + 2)log10 = log(1632) Divide both sides of equation by log10. x + 2 = 3.21 Subtract 2 on both sides of equation. x = 1.21 If you found this helpful, I encourage to subscribe to our youtube channel https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)